OpenStudy (anonymous):

Reducing and multiplication with fractions i seriously need help with this there is a time limit pleeeeeease 18/54,11/25,50/9.find the product in lowest terms

OpenStudy (anonymous):

\[\frac{22}{27} \]

OpenStudy (anonymous):

Take the product of the numerators and divide the result by the product of the denominators. Simplify the result.

OpenStudy (anonymous):

<333

OpenStudy (anonymous):

even if they dont have the same denominators?

OpenStudy (anonymous):

\[\frac{9900}{12150}=\frac{22}{27} \]

OpenStudy (anonymous):

\[\frac{18\ 11\ 50}{54\ 25\ 9} \]